write a formula to solve:
you have $6.70 in quarters and dimes, a total of 46 coins, how many quarters and how many dimes

Answer by jojo14344(1513)

Just a formula?.. SURE!
Let "x"= # of Quarters
"y"= #of dimes
So, x+y=46 coins right? Mark this equation 1.
Next, 0.25(quarter)*x+0.10(dime)*y=$6.70 right? Mark this equation 2.
In eqn 1 you get "y=46-x" and substitute this value of y in equation 2.
Therefore your working formula will be:
0.25(x)+0.10(46-x)=$6.70
x=14, # of quarters
for y, go back eqn 1, y=46-14
y=32, # of dimes
In doubt? Go back eqn 2 to check,
0.25(14)+0.10(32)=$6.70
$3.50+$3.20=$6.70
$6.70=$6.70 cool!
Sorry amigo, you're just asking for the formula but I got carried away solving it all the way.
Thank you,
Jojo